What is the slope and y-intercept of the equation 3y=4x+1
Answer by tutorcecilia(2144)

3y=4x+1 [Solve for y so the equation is in the slope-intecept format]
y=4x/3 + 1/3 [Simplify]
y = 4/3x + 1/3 [Slope-intercept format]
y = m x + b
Slope = m = 4/3
y-intercept = b = 1/3